RawlsK29:

What is the value of a in the equation 4a + 3(a + 1) = 24? 2 3 7 21

RawlsK29:

@563blackghost @Elsa213 @ThisGirlPretty

dude:

Ah we want to isolate a, do you have an idea on what to do first?

RawlsK29:

4a + 3(a) + 3(1) = 24?

RawlsK29:

not sure though?

dude:

Yes that is right

RawlsK29:

4a + 3a + 3 = 24 7a - 3 + 3 7a = 24 is this right so far?

dude:

Almost \(4a + 3a + 3 = 24\) \(7a - 3 + 3\) <-When subtracting make sure to do it to both sides of the equation So... \(7a +3\color{red}{- 3}=24 \color{red}{- 3}\\ 7a=21 \)

RawlsK29:

Ohhh i get it now, so 7a = 21 7a/7 and 21/7 a = 3

BlankSpace:

Step 1: Simplify both sides of the equation. 4a+3(a+1)=24 4a+(3)(a)+(3)(1)=24(Distribute) 4a+3a+3=24 (4a+3a)+(3)=24(Combine Like Terms) 7a+3=24 7a+3=24

BlankSpace:

Step 2: Subtract 3 from both sides. 7a+3−3=24−3 7a=21

dude:

Correct it is 3

RawlsK29:

Thanks so much!(:

BlankSpace:

Step 3: Divide both sides by 7. 7a/7 = 21/7 a=3

BlankSpace:

So yeah, 3.

dude:

If you want to check you can substitute 3 back for a in the original equation \(\checkmark Check\\4(3) + 3((3) + 1) = 24\\ 12+3(4)=24\\ 12+12=24\\ \color{green}{\checkmark 24=24}\)